OpenStudy (anonymous):

3tan^2*X+cot^2*X=4 X= ??

OpenStudy (ranga):

Make the substitution cot(X) = 1/tan(X), multiply throughout by tan^2(X). Let y = tan^2(X). You will get a quadratic equation in y. Solve for y and then put back y = tan^2(X) and solve for X.

OpenStudy (anonymous):

you can show me more? im stuck

OpenStudy (ranga):

3tan^2*X + cot^2*X = 4 cot(X) = 1/tan(X) 3tan^2(X)+1/tan^2(X) = 4 Multiply throughout by tan^2(X) 3tan^4(X) + 1 = 4tan^2(X) 3tan^4(X) - 4tan^2(X) + 1 = 0 Let y = tan^2(X) 3y^2 - 4y + 1 = 0 Solve the quadratic equation for y.

OpenStudy (anonymous):

thanks man!

OpenStudy (ranga):

After you solve for y, remember to replace y by tan^2(X) and solve for X. You are welcome.